Question 8. (15 points total) (Calculating financial ratios) Use the balance sheet and income statement for the J. P. Robard Mfg. Company to calculate the following ratios: Current ratio (Round to two decimal places ) Times interest earned (Round to two decimal places.) Inventory turnover (Round to two decimal places.) Total asset turnover (Round to two decimal places ) Operating profit margin (Round to one decimal places.) Operating return on assets (Round to one decimal places,) Debt ratio (Round to one decimal places.) times times Average collection period (Round to one decimal places) days Fixed asset turnover (Round to two decimal places.) Return on equity (Round to one decimal places.) J. P. Robard Mfg., Inc. Balance Sheet ($000) Cash Accounts receivable Inventories $460 1,920 910 3,290 4,480 S7,770 Current assets Net fixed assets Total assets Accounts payable Accrued expenses Shoet-term notes payable $1.170 570 260 $2,000 1,920 3,850 Current Eabilities Long-term debt Owners equity Total lEablities and owners equity J. P. Robard Mfg., Inc. Income Statement (S000) Net sales (all credit) Cost of goods sold $7,990 (3,340) 4,650 (3,000) $1,650 (363) $1,287 (515) $772 Gross profit Operating expenses (inchudes $500 depreciation) Net operating income Interest expense Earnings before taxes Income taxes (40%) Net income
Answer #1
Solution:
Current ratio 1.65 times
Times interest earned 4.55 times
Inventory turnover 3.67 times
Total asset turnover 1.03 times
Operating profit margin 20.7 %
Operating return on assets 21.2 %
Debt ratio 50.5 %
Average collection period 87.7 days
Fixed asset turnover 1.78 times
Return on equity 20.1 %
Working Notes:
Current ratio = Current Assets/Current Liabilities
Current ratio = 3,290/2000
Current ratio = 1.645
Current ratio = 1.65
Times interest earned
Times Interest Earned = EBIT/Interest Expense
Times Interest Earned = operating income /Interest Expense
Times Interest Earned =$1,650 /$363
Times Interest Earned =4.545454545
Times Interest Earned =4.55
Inventory turnover
Inventory Turnover = COGS/Inventory
Inventory Turnover = 3340/910
Inventory Turnover = 3.67032967
Inventory Turnover = 3.67
Total asset turnover
Total Asset Turnover = Sales/Total Assets
Total Asset Turnover =7,990/7,770
Total Asset Turnover =1.028314028
Total Asset Turnover =1.03
Operating profit margin
Operating Profit Margin = Operating Earnings/Revenue
Operating Profit Margin = $1,650/$7,990
Operating Profit Margin = 0.206508135
Operating Profit Margin = 20.7%
Operating return on assets
Operating Return on Assets = EBIT/Average total assets
Operating Return on Assets =operating income /Average total assets
Operating Return on Assets =$1,650 /7,770
Operating Return on Assets =0.212355212
Operating Return on Assets =21.2%
Debt ratio
Debt Ratio = Total Debt / Total Assets
Debt Ratio =(Current liabilities + Long term debt)/ Total Assets
Debt Ratio =($1,920 + $2,000)/ $7,770
Debt Ratio =0.504504505
Debt Ratio =50.5%
Average collection period
Average Collection Period = (Average Account Receivable)/(Annual Sales/365)
Average Collection Period = (1920)/(7990/365)
Average Collection Period = 87.70963705
Average Collection Period = 87.7 days
Fixed asset turnover
Fixed asset turnover = Net Sales/Average fixed assets
Fixed asset turnover = 7990/4,480
Fixed asset turnover = 1.783482143
Fixed asset turnover = 1.78
Return on equity
Return to Equity = Net Income/Shareholder’s Equity
Return to Equity = 772/3,850
Return to Equity = 0.200519481
Return to Equity = 20.1%
Please feel free to ask if anything about above solution in comment section of the question.